OpenStudy (anonymous):

whats the DENOMINATOR of 2√(5^3)/6+3√(10^5) a.2√(5^3)/3√(10^5) b. 2√(5^3) c. 3√(10^5) d. 6+3√(10^5)

OpenStudy (mertsj):

The denominator is the bottom of the fraction.

OpenStudy (anonymous):

I suppose you mean \[\frac{2\sqrt{5^3}}{6+3\sqrt{10^5}}\], so d.